# Cardinality of Real Numbers and Set Axiomatization

### v2

@@ 2 Discussion
Let us first follow the Löwenheim-Skolem theorem and expand ZF model with AC. While AC is independent of the ZF set theory, which means that both the axiom itself and its negation are consistent with ZF, AC expands ZF. It allows the construction of set $S^{\prime}$, which is not possible without AC, and to recognize in the set $S^{\prime}$ every possible Cauchy sequence representing any number, even if no sequence is ever explicitly defined. It also allows the construction of any sequences of any number from the intervals with algebraic endpoints that are present in the set, similarly to some kind of a variation of the ‘$V=L$’ approach. To demonstrate this explicitly, in example 1, a countable set $S^{\prime}$ will be considered, which contains all pairs of algebraic numbers and transcendental numbers generated by a function $f$ that places an arbitrarily selected transcendental number between each of these pairs of algebraic numbers. Then using AC, it will be shown that such a set is equivalent to the set of real numbers.
Example 1. Consider a closed interval $[a,b]$, where $a<b$ are both algebraic numbers. We define the set $S$, which consists of all possible closed intervals $I_{n}^{\prime}=[a_{n},b_{n}]\cap\overline{\mathbb{Q}}$ within $[a,b]$, with the specification that $I_{n}^{\prime}$ are considered in the context of containing only algebraic numbers, i.e., they are the intersection of the intervals $[a_{n},b_{n}]$ with the set of all algebraic numbers, denoted as $\overline{\mathbb{Q}}$, and will be denoted as $[a_{n},b_{n}]^{\prime}$. Here, $a_{n}$ and $b_{n}$ are algebraic numbers, and $a_{n}\leqslant b_{n}$. Thus, the set $S$ encompasses every such algebraic interval $I_{n}^{\prime}=[a_{n},b_{n}]\cap\overline{\mathbb{Q}}=[a_{n},b_{n}]^{\prime}\subseteq[a,b]$.
Next, we define a function $f$ that, for each interval $I_{n}^{\prime}$, selects an arbitrarily chosen transcendental number $t_{n}$, ensuring that $t_{n}$ falls within the algebraic bounds of $I_{n}^{\prime}$, i.e., $a_{n}\leqslant t_{n}\leqslant b_{n}$, and the function $f$ has to act on all intervals $I_{n}^{\prime}$ in the set $S$. This setup positions $t_{n}$ as associated with the interval $I_{n}^{\prime}$, conceptualizing the inclusion of $t_{n}$ in the space spanned by $I_{n}^{\prime}$, despite $t_{n}$ not being an algebraic number. Formally, the function $f$ is described as follows:
$f:S\rightarrow\mathbb{R}\setminus\overline{\mathbb{Q}},\mbox{ such that for all }I_{n}^{\prime},a_{n}\leqslant f(I_{n}^{\prime})\leqslant b_{n},\mbox{ and}f(I_{n}^{\prime})=t_{n}$ (1)
In addition, it is required that the function $f$ also satisfies: a) that it acts on all intervals $I_{n}^{\prime}$ simultaneously, not consecutively. There is no difference in outcomes whether the function is acting on intervals successively or adding numbers simultaneously; the outcome must be the same since it depends only on the properties of intervals. However, this requirement is to avoid discussions such as whether the process of inserting numbers can be completed, and to ensure that function $f$ is not only acting on individual intervals $I_{n}^{\prime}$, but also on the entire sequences of such intervals if such sequences exist in the set $S$. Thus, if in the set $S$ there are explicitly or implicitly defined sequences of nested intervals $I_{1}^{\prime}\supset I_{2}^{\prime}\supset I_{3}^{\prime}\supset\ldots$, when function $f$ is inserting numbers $t_{n}$ into each of these intervals, creating intervals $I_{n}^{\prime\prime}$, which in addition to algebraic numbers also contain numbers $t_{n}$, then the generated numbers $t_{n}$ must also satisfy the condition that there will be a number $t$ that fits into all such intervals when the intervals are nested and function $f$ by definition acted on all of them. Therefore the condition to insert numbers in all intervals at same time implies confrontation of the function $f$ with all intervals, with the entire sequence, which requires that there must be a number that is in the intersection of all of them since all intervals are overlapping and each interval $I_{n}^{\prime\prime}$ is nonempty. Therefore, when confronted with infinite sequences the function $f$ will also have to generate numbers $t_{n}$ that satisfy
$I_{1}^{\prime\prime}\supset I_{2}^{\prime\prime}\supset I_{3}^{\prime\prime}\supset...=\bigcap_{n\in\mathbb{N}}\{I_{n}^{\prime\prime}\}=\{t\}$ (2)
b) In addition, it is required that if a transcendental number generated by the function $f$ is in an interval $I_{n}^{\prime\prime}$, then it must also be contained in all superintervals that contain that interval, and c) that the function acts only on the intervals $I_{n}^{\prime}$ that do not contain transcendental numbers and always generates a different transcendental number $t_{n}$. Additionally, the intervals $I_{n}^{\prime}=[a_{n},b_{n}]\cap\overline{\mathbb{Q}}$ include all subintervals $[a_{m},b_{m}]\cap\overline{\mathbb{Q}}$ contained within $[a,b]$.
Let us denote such a created set of everywhere dense transcendental numbers $t_{n}$ as $T$, i.e.
$T=\{f([a_{n},b_{n}]\cap\overline{\mathbb{Q}}):\mbox{ for all }[a_{n},b_{n}]\cap\overline{\mathbb{Q}}\in[a,b]\}=\bigcup_{n\in\mathbb{N}}\{t_{n}\}$ (3)
Adding set $T$ to set $S$ creates set $S^{\prime}$ that contains all algebraic numbers, all intervals $I_{n}^{\prime}$ with algebraic endpoints, and also all transcendental numbers $t_{n}$ from set $T$, which are inserted by function $f$ in the $I_{n}^{\prime}$ intervals.
$S^{\prime}=\bigcup_{n\in\mathbb{N}}\{a_{n},t_{n},b_{n}\}=\overline{\mathbb{Q}}\cup T$ (4)
The cardinality of set $S^{\prime}$ and its relation to the cardinality of real numbers can then be formulated through the following theorem:
Theorem 1: Consider a closed interval $[a,b]$, where $a<b$ are algebraic numbers. Define a set $S$ to consist of all closed subintervals $I_{n}^{\prime}=[a_{n},b_{n}]\cap\overline{\mathbb{Q}}$ within $[a,b]$, where $a_{n}$ and $b_{n}$ are algebraic endpoints satisfying $a_{n}\leqslant b_{n}$, and each $I_{n}^{\prime}$ is populated exclusively by algebraic numbers. Concurrently, define the set $T$ to encompass only transcendental numbers $t_{n}$, produced by the function $f:S\rightarrow\mathbb{R}\setminus\overline{\mathbb{Q}}$, which assigns a unique transcendental number $t_{n}=f(I_{n}^{\prime})$ to each interval $I_{n}^{\prime}$, in such a manner that $a_{n}\leqslant t_{n}\leqslant b_{n}$ holds for all elements of $S$, and that the function $f$ has to act simultaneously on all elements, all $I_{n}^{\prime}$, in the set $S$. This operation by $f$ effectively simultaneously augments each interval $I_{n}^{\prime}$ with a transcendental number $t_{n}$, yielding enhanced intervals $I_{n}^{\prime\prime}$, which can then be conceptualized as containing both the original algebraic numbers and the newly added transcendental numbers $t_{n}$. Consequently, the augmented set $S$, denoted as $S^{\prime}=S\cup T$, is both denumerable and complete.
A simplified version of this theorem states: Construct a set $S$ that includes all pairs of algebraic numbers $\{a_{n},b_{n}\}$ and simultaneously insert a single transcendental number $t_{n}$ between each pair, such that $t_{n}=f(\{a_{n},b_{n}\})$ and $a_{n}\leqslant t_{n}\leqslant b_{n}$ for every pair, and do this for all pairs in $S$. This results in a new set $S^{\prime}$ that is both complete and countable.
Proof of Theorem 1. The proof is based on a counterexample. Let us assume that there is a missing transcendental number $T_{x}$ that is different from all $t_{n}$ numbers generated by the function $f$ and therefore from all numbers in the set $S^{\prime}$. It must be a transcendental number since all algebraic numbers are, by definition, included in the set $S^{\prime}$. Then define that missing number $T_{x}$ with a sequence of nested closed intervals $[\alpha_{n},\beta_{n}]$ with algebraic endpoints $\alpha_{n}$ and $\beta_{n}$, such that $T_{x}$ is defined by that sequence
$[\alpha_{1},\beta_{1}]\supset[\alpha_{2},\beta_{2}]\supset[\alpha_{3},\beta_{3}]…\supset[\alpha_{n},\beta_{n}]...$ (5)
and it is the intersection of these intervals
$T_{x}=\bigcap_{n\in\mathbb{N}}[\alpha_{n},\beta_{n}].$ (6)
Using the AC, one can identify in the set $S^{\prime}$ closed intervals $I_{n}^{\prime\prime}=[a_{n},b_{n}]\cap\overline{\mathbb{Q}}\cup T$ that will be denoted $[a_{n},b_{n}]^{\prime\prime}$ that have the same endpoints as the closed intervals $[\alpha_{n},\beta_{n}]$ in the sequence (5), i.e., closed intervals $I_{n}^{\prime\prime}=[a_{n},b_{n}]^{\prime\prime}$ with $a_{n}=\alpha_{n}$ and $b_{n}=\beta_{n}$. Every closed interval in the sequence (5) can be recognized in the set $S^{\prime}$ because the set $S^{\prime}$ contains all closed intervals with algebraic endpoints, and therefore all intervals that are in the sequence (5). All these closed intervals $I_{n}^{\prime\prime}$ can be recognized by AC and assembled together to create a sequence that also represents the number $T_{x}$, i.e., the sequence of $I_{n}^{\prime\prime}$ intervals
$[a_{1},b_{1}]^{\prime\prime}\supset[a_{2},b_{2}]^{\prime\prime}\supset[a_{3},b_{3}]^{\prime\prime}…\supset[a_{n},b_{n}]^{\prime\prime}...$ (7)
and $T_{x}$ is the intersection of these $I_{n}^{\prime\prime}$ intervals
$T_{x}=\bigcap_{n\in\mathbb{N}}I_{n}^{\prime\prime}.$ (8)
So, using the AC, one can recognize in the set $S^{\prime}$ the sequence (7) that is exactly the same as sequence (5), with closed intervals $I_{n}^{\prime\prime}$ having the same endpoints as intervals in sequence (5). To demonstrate that the number $T_{x}$ is not missing, that it is contained within the set $S^{\prime}$, it is sufficient to show that the intersection (8) of the closed intervals $I_{n}^{\prime\prime}$ in sequence (7) is not empty and that it contains the number $T_{x}$. To proceed further, let us first prove the following lemma.
Lemma 1. Let $T_{y}$ be a real number defined as the limit of an infinite converging nested sequence of closed intervals $I_{n}=[\alpha_{n},\beta_{n}]$, satisfying the following conditions:
  1. 1.
    $I_{n+1}\subseteq I_{n}$ for all $n$,
  2. 2.
    $|\beta_{n}-\alpha_{n}|\to 0$ as $n\to\infty$, and
  3. 3.
    $\alpha_{n}$ and $\beta_{n}$ are algebraic for all $n$.
The requirement to select a single number from each interval $I_{n}$, and to continue this selection throughout all sequence intervals, implies that it is impossible to avoid choosing the number $T_{y}$. This inevitability arises because, as $n$ approaches infinity, all these overlapping intervals collectively represent the single number $T_{y}$ on the real line.
Proof of Lemma 1. In a sequence of nested closed intervals, each chosen number at step $n$ must reside within the intersection of all prior intervals, reflecting the sequence’s nested structure. As the sequence progresses, the intersection narrows, ultimately resulting in a single point, $T_{y}$, as stipulated by the NIP. This convergence limits selection options exclusively to $T_{y}$ when traversing the entire sequence. Therefore, it is impossible to avoid selecting $T_{y}$, as it becomes the sole viable number within the intersection of all intervals, given that each selected number must be from the intersection of intervals, which eventually contains only the single number $T_{y}$. This outcome is evident when considering that the overlapping intervals collectively represent $T_{y}$.

### v3

@@ 2 Discussion
Let us first follow the Löwenheim-Skolem theorem and expand ZF model with AC. While AC is independent of the ZF set theory, which means that both the axiom itself and its negation are consistent with ZF, AC expands ZF. It allows the construction of set $S_{m}$, which is some kind of a variation of the ‘$V=L$’ approach. As described in Theorem 1, set $S_{m}$ is generated by marking or identifying an arbitrarily selected irrational number within each interval with rational endpoints. It will be shown that such a countable set is everywhere dense in $\mathbb{R}$, and Dedekind complete, i.e., that it has properties of set $\mathbb{R}$.
2.1 Main Construction and Theorem
Let us start with a set of rational numbers and, in a constructive way under the ZFC formalism, create a set equivalent to the set of real numbers.
Theorem 1 (Construction and properties of the set $S_{m}$).
Enumerate all closed intervals with rational endpoints:
$S_{Q}=\{I_{n}=[a_{n},b_{n}]:a_{n},b_{n}\in\mathbb{Q},\;a_{n}\leq b_{n},\;n\in\mathbb{N}\}.$ (2)
Suppose each interval $I_{n}$ is assigned exactly one irrational mark $x_{n}\in I_{n}$ under the following structural marking axioms:
  1. (S1)
    Coverage: Every interval $I_{n}$ receives exactly one irrational mark:
    $\forall n\in\mathbb{N},\quad x_{n}\in I_{n}.$ (3)
  2. (S2)
    Nested Interval Marking: Every subinterval with rational endpoints of a marked interval, $[a^{\prime},b^{\prime}]\subseteq I_{n}$, is also marked from the outset. Thus, no interval or subinterval, regardless of its depth or nesting, remains unmarked. This implies that each $I_{n}$ contains infinitely many marked points.
  3. (S3)
    Global Simultaneous Marking: All intervals and subintervals are marked simultaneously in a global, structurally coherent manner. Specifically, the marking does not depend on sequential adaptation but is preassigned once and for all, fully respecting the nested and overlapping structure of all its subintervals.
  4. (S4)
    Universal Markability: For every interval $I_{n}=[a_{n},b_{n}]$, any point $x_{n}\in I_{n}$ is eligible to be assigned as a mark $x_{n}\in M$.
Let the set of all such marks be:
$M=\{x_{n}:n\in\mathbb{N}\}\subset\mathbb{R}\setminus\mathbb{Q}.$ (4)
and structurally marked set
$S_{m}:=\mathbb{Q}\cup\{x_{n}:n\in\mathbb{N}\}=\mathbb{Q}\cup M.$ (5)
Then the set $S_{m}$ is countable, everywhere dense in $\mathbb{R}$, and Dedekind complete, that is, every Dedekind cut in $S_{m}$ is realized by an element of $S_{m}$.
Proof:
Countability
$\mathbb{Q}$ and $M$ are both countable, so $S_{m}$ is a countable union of countable sets.
Density
Since $\mathbb{Q}\subset S_{m}$ and $\mathbb{Q}$ is dense in $\mathbb{R}$, it follows that $S_{m}$ is dense in $\mathbb{R}$.
Dedekind Completeness
To prove Dedekind’s completeness, let us first consider this lemma:
Lemma 1 (Dedekind Completeness of $S_{m}$ via Gap‐Point & Squeeze).
Under the structural marking axioms (S1)–(S4) of Theorem 1, the set $S_{m}=\mathbb{Q}\cup M$ is Dedekind complete: every Dedekind cut in $S_{m}$ is realized by an element of $S_{m}$.
Proof.
Let $A\mid B$ be a Dedekind cut in $S_{m}$, and set
$t=\sup(A\cap\mathbb{Q})=\inf(B\cap\mathbb{Q})\in\mathbb{R}.$ (6)
We show $t\in S_{m}$.
Case 1: $t\in\mathbb{Q}$. Then $t\in\mathbb{Q}\subset S_{m}$, so the cut is filled.
Case 2: $t\notin\mathbb{Q}$. We must prove $t\in M$.
(a) Nested‐interval representation. Since $t$ is the gap‐point in $\mathbb{Q}$, choose
$a_{1}<a_{2}<\cdots<t<\cdots<b_{2}<b_{1},\quad b_{n}-a_{n}<\tfrac{1}{n},$ (7)
and let $I_{n}=[a_{n},b_{n}]$. Then
$I_{1}\supseteq I_{2}\supseteq\cdots,\quad\text{and}\quad\bigcap_{n=1}^{\infty}I_{n}=\{t\}.$ (8)
(b) Define $\alpha,\beta$. Let
$\alpha=\sup\{x\in M:x<t\},\qquad\beta=\inf\{x\in M:x>t\}.$ (9)
Suppose that $t\notin M$, we only know $\alpha\leq t\leq\beta$.
Subcase 2a (strict gap $\alpha<\beta$). Pick rationals $r^{-}<t<r^{+}$ with $\alpha<r^{-}<t<r^{+}<\beta$. For large $n$,
$I_{n}\subset(r^{-},r^{+}),$ (10)
and by definition no mark lies in $(r^{-},r^{+})$, so $I_{n}$ would be unmarked—contradicting (S1). Thus $\alpha<\beta$ is impossible.
Subcase 2b (cluster $\alpha=\beta=t$). Here every mark in $M$ accumulates at $t$. Now:
1. By (S1)–(S3), each $I_{n}$ is marked, so pick $x_{n}\in M\cap I_{n}$. Thus, for each $n$, both $x_{n}\in M$ and $t\in I_{n}$.
2. Since $\mathrm{diam}(I_{n})\to 0$ and $t\in I_{n}$, we have $|x_{n}-t|\leq\mathrm{diam}(I_{n})\to 0$, hence $x_{n}\to t$.
3. For any fixed $x_{k}\neq t$, set $\delta_{k}=\tfrac{1}{2}|x_{k}-t|>0$. Then for all $n$ with $\mathrm{diam}(I_{n})<\delta_{k}$, we have $I_{n}\subset(t-\delta_{k},t+\delta_{k})$, so $x_{k}\notin I_{n}$. Hence no $x_{k}\neq t$ can lie in every $I_{n}$. The Nested Interval Property ensures that all $x_{k}\neq t$ are eventually excluded, and only $t$ survives.
4. Yet each $I_{n}$ must contain some mark. Axiom (S2) requires that there are marked points $x\in M$ in every subinterval that, together with $t$, remain in every subinterval, regardless of how small, all the way to the intersection. If no mark survives through all intervals together with $t$, some deep subintervals would be left unmarked, contradicting (S2).
Thus, the only point able to survive all levels is $t$ itself. Therefore, one of the $x_{n}$ is forced by the structure of the intervals not only to converge to $t$ but to be equal to $t$, which is allowed by (S4), and hence $t\in M$.
In both subcases, we reach the conclusion that $t\in M\subset S_{m}$, contradicting the assumption that $t\notin S_{m}$. Thus, every Dedekind cut in $S_{m}$ is filled, and $S_{m}$ is Dedekind complete. ∎
Conclusion of Theorem 1
By combining the results, it follows that $S_{m}$ satisfies all the stated properties. Specifically, $S_{m}$ is everywhere dense in $\mathbb{R}$, Dedekind complete, and countable. Therefore, $S_{m}$ represents a rigorous construction of a set that is countable yet exhibits the properties of the real numbers. This proves the main Theorem 1. $\square$
This outcome is unexpected, but not comppletely, since for any arbitrarily selected numbr $r$ we can find in the set $S_{Q}$ collection of intervals $[a_{n},b_{n}]$, every contining marked numbers, that are closing on $r$ and after an infinite overlapping will include only one marked numer that must be $r$.
To demonstrate that select a number $r$ and recognize that the set $S_{Q}$ consisting of intervals with rational endpoints is dense in $\mathbb{R}$, for any $\epsilon>0$, there exist intervals $[a_{n},b_{n}]\in S_{Q}$ that can be iteratively selected with smaller $\epsilon_{n}\to 0$ as $n\to\infty$, constructing a sequence of nested intervals $[a_{n},b_{n}]$ satisfying:
$r-\epsilon\leqslant a_{n}\leqslant r\leqslant b_{n}\leqslant r+\epsilon.$ (11)
Therefore in the set $S_{Q}$ it is always possible to recognize and select an infinite sequence of nested intervals $S_{Q_{n}}$, or better say in set $S_{Q}$ there exists an implicit, infinite sequence of nested nonempty intervals $S_{Q_{n}}$ with endpoints that close on $r$ and uniquely select the point $r$, as:
$a_{n}\leqslant r\leqslant b_{n},\mbox{ for all }n.$ (12)